OpenStudy (mwhit2101):

Need a quick answer to this equation please: 3x-4+2=8

OpenStudy (studytheworld):

Hi! x=10/3. Do you know how we got this answer?

TheSmartOne (thesmartone):

3x - 4 + 2 = 8 First simplify what is -4 + 2 = ? Hint: It's the same thing as 2 - 4 = ?

OpenStudy (mwhit2101):

-2